OpenStudy (accidentalaichan):

What is the simplified form of sqrt r^49?

OpenStudy (accidentalaichan):

\[\sqrt{r ^{49}}\]

OpenStudy (anonymous):

7

OpenStudy (jdoe0001):

heheh

OpenStudy (accidentalaichan):

\[r ^{7}\]?

OpenStudy (jdoe0001):

how many times is "2" into 49?

OpenStudy (jdoe0001):

49 is an odd number, thus it'd end up with some remainder of 1 but the amount of "2" into ... will be?

OpenStudy (accidentalaichan):

I know it's 7, but this is why I'm confused.

OpenStudy (jdoe0001):

hmm how do you know is 7 anyway?

OpenStudy (accidentalaichan):

There has to be two rational numbers that create the square. Otherwise, it's not a perfect square.

OpenStudy (jdoe0001):

well.... so you're saying that \(\bf r^{49} = 7\) whatever happened to the "r"? it sorta went "missing in action"

OpenStudy (jdoe0001):

ahemm mehh anyhow \(\bf \sqrt{r^{49}} = 7\) rather

OpenStudy (accidentalaichan):

It'd have to be \[7r \sqrt{r} \], because you'd multiply 7r by r and get r^49

OpenStudy (accidentalaichan):

At least that's what the lesson says

OpenStudy (jdoe0001):

well. that'd be true for 49r.... even then.. the "r" is sorta missing

OpenStudy (jdoe0001):

how many times does "2" go into 49 you think?

OpenStudy (accidentalaichan):

24.5

OpenStudy (jdoe0001):

24.5 yes.... well. if we throw away the .5 24 times so 24 * 2 = 48 so \(\Large \bf { \sqrt{r^{49}}\to \sqrt{r^{48+1}}\to \sqrt{r^{24\cdot 2+1}}\to \sqrt[{\color{red}{ 2}}]{(r^{24})^2\cdot r} }\)

OpenStudy (accidentalaichan):

So, I'm guessing it is \[r ^{24}\sqrt{r}\] ?

OpenStudy (dan815):